OpenStudy (anonymous):

write equation in slope form y=5/8X -1 X-3/4Y=1/2 3X-13=2Y X=3/5-2/3Y Y=2X-6

OpenStudy (anonymous):

some of them are in slope form already - like y = 2x - 6 y = 5/8 x - 1 second one 3/4 y = x - 1/2 divide thru by 3/4 y = 4/3 x - 1/3 third y = 3/2 x - 13/2 4th 2/3 y = -x + 3/5 y = -3/2 x + 3/10

OpenStudy (anonymous):

ok?